Workers: Akutsu, Michimura, K. Tanaka (17:30-21:30)
GPT part installation
Installed optics for green path from GBS1, except for green QPD boards (See JGW-T1808962 for optical layout)
- we marked screw holes on optical table every 10 holes
- we labeled all the mirror mounts with the part number of the optic
- the position of GST2 was 25mm too far from GLNS1, so we moved it to the default position
- beam seems to be focused at around GST3, as expected
- Other things
- Height of Gig-E camera is a bit short
- All the LVDT outputs went close to zero from around 21:26. We didn't notice this until we got back to the control room.
- Align and balance the BRT and the TMS-VIS.
- Detach the remaining white tapes just before closing the chamber; be noted that there are white tapes on the PEEK connectors at the inside of the flanges.
- Torque completely the bolts which are fixing the upper and the lower base plates.
- Insert metal gaskets before torquing the ICF203 flanges with Dsub connectors.
- Modify MEDM screen
- Add DC PD and so on.
Cable Dsub 15pin for TMS-VIS OpLev.
- Modify cabling for DC PD.
- Return plastic box.
- Install rails in digital rack.
- They may be necessary under LVDT distributor, geophone distributor, and WHT board.
- Nagano is not sure it is necessary under HP coil driver.
- Consider how to cable Dsub cables to QPD.
- Investigate LVDT outputs
- Install QPD boards and do the final cabling
- Check why Gig-E camera is not working
- Install IR optics
- M4-12mm (or longer) screws instead of M4-10mm for fixing mirror holders to 20mm dia. posts
- Spacer for Gig-E camera (height is shorter than 4-inch right now)